### v2

# Geometric Analogues of Mie Scattering

Abstract
Cross-sections for particles scattered from selected spatial geometries exhibit many of the same interesting features as Mie scattering.
1 Introduction
Scattering of light from dielectric and/or conductive spheres (so-called Lorenz-Mie scattering) is well-known to exhibit fascinating but very complicated structure [1].  There are numerous poles in the various scattering amplitudes, resulting in very narrow as well as rather broad resonances, and for resistive spheres there can be substantial power loss, i.e. inelastic scattering.
Similar effects are easily produced by spherical scattering centers consisting of nontrivial spatial geometries and nothing else [2, 3, 4].  A detailed discussion of selected examples involving such alternative scattering mechanisms is the subject of this paper.  The analysis is carried out in the context of the three-dimensional Helmholtz equation wherein only spatial geometry is modified from Euclidean space.  Time is taken to be universal to all frames.  In this sense the study is non-relativistic.
There are known equivalence relations for scattering by potentials, by dielectric materials, and by various geometric features [5, 6].  Such equivalences are not invoked here.  Rather, the emphasis here is on the geometry and the resulting cross-sections as computed directly from partial wave expansions.  Nevertheless, as a practical matter, the analysis to follow may provide insight into the behavior of mathematically equivalent meta-materials.  Alternatively, the following may provide useful intuition about scattering phenomena in general relativity, at least in the non-relativistic limit.  There are some previous investigations along these same lines [7, 8], including studies of the corresponding Born series, which place emphasis on other aspects of the same subject.
Relevant formulae are given in the body of the paper along with selected numerical graphs to illustrate various important physical effects.  Additional plots are collected in an Appendix.
2 Scattering by spheres
A brief review of scattering by a homogeneous sphere of radius $R$ embedded in three-dimensional Euclidean space is warranted for purposes of comparison.  Scattering of plane waves by impenetrable “hard” spheres in non-relativistic quantum mechanics and by perfectly conducting spheres in classical electromagnetic theory are described by the following integrated cross-sections [9, 10, 11].
$\sigma_{\text{QM hard sphere}}=\sum_{l=0}^{\infty}\sigma_{l\text{ QM hard sphere}}\ ,\ \ \ \sigma_{l\text{ QM hard sphere}}=\frac{4\pi}{k^{2}}\left(2l+1\right)\sin^{2}\left(\delta_{l}\right)$ (1)
$\sigma_{\text{EM perf cond sphere}}=\sigma_{\text{TE perf cond sphere}}+\sigma_{\text{TM perf cond sphere}}$ (2)
$\sigma_{\text{TE perf cond sphere}}$ $=\sum_{l=1}^{\infty}\sigma_{l\text{ TE perf}}\ ,\ \ \ \sigma_{l\text{ TE perf}}=\frac{2\pi}{k^{2}}\left(2l+1\right)\sin^{2}\left(\delta_{l}\right)$ (3)
$\sigma_{\text{TM perf cond sphere}}$ $=\sum_{l=1}^{\infty}\sigma_{l\text{ TM perf}}\ ,\ \ \ \sigma_{l\text{ TM perf}}=\frac{2\pi}{k^{2}}\left(2l+1\right)\sin^{2}\left(\Delta_{l}\right)$ (4)
$e^{2i\delta_{l}}=-h_{l}^{\left(2\right)}\left(kR\right)/h_{l}^{\left(1\right)}\left(kR\right)\ ,\ \ \ e^{2i\Delta_{l}}=-\frac{d}{dR}\left(Rh_{l}^{\left(2\right)}\left(kR\right)\right)/\frac{d}{dR}\left(Rh_{l}^{\left(1\right)}\left(kR\right)\right)$ (5)
The scattering is purely elastic with real phase shifts $\delta_{l}$ and $\Delta_{l}$.  With the use of the Bessel function relations
$h_{l}^{\left(1,2\right)}\left(x\right)$ $=\sqrt{\frac{\pi}{2x}}\left(J_{l+\frac{1}{2}}\left(x\right)\pm iY_{l+\frac{1}{2}}\left(x\right)\right)$ (6)
$\frac{d}{dx}\left(\sqrt{x}J_{l+1/2}\left(x\right)\right)$ $=\frac{1}{\sqrt{x}}\left(\left(l+1\right)J_{l+\frac{1}{2}}\left(x\right)-xJ_{l+\frac{3}{2}}\left(x\right)\right)$ (7)
$\frac{d}{dx}\left(\sqrt{x}Y_{l+1/2}\left(x\right)\right)$ $=\frac{1}{\sqrt{x}}\left(\left(l+1\right)Y_{l+\frac{1}{2}}\left(x\right)-xY_{l+\frac{3}{2}}\left(x\right)\right)$ (8)
it follows that
$\frac{1}{\pi R^{2}}~\sigma_{\text{QM hard sphere}}=\frac{4}{x^{2}}\sum_{l=0}^{\infty}\left(2l+1\right)\frac{\left(J_{l+1/2}\left(x\right)\right)^{2}}{\left(J_{l+1/2}\left(x\right)\right)^{2}+\left(Y_{l+1/2}\left(x\right)\right)^{2}}$ (9)
$\frac{1}{\pi R^{2}}~\sigma_{\text{TE perf cond sphere}}=\frac{2}{x^{2}}\sum_{l=1}^{\infty}\left(2l+1\right)\frac{\left(J_{l+1/2}\left(x\right)\right)^{2}}{\left(J_{l+1/2}\left(x\right)\right)^{2}+\left(Y_{l+1/2}\left(x\right)\right)^{2}}=\frac{1}{2}~\frac{\sigma_{\text{QM hard sphere}}}{\pi R^{2}}$ (10)
$\frac{1}{\pi R^{2}}~\sigma_{\text{TM perf cond sphere}}=\frac{2}{x^{2}}\sum_{l=1}^{\infty}\left(2l+1\right)\frac{\left(\left(l+1\right)J_{l+\frac{1}{2}}\left(x\right)-xJ_{l+\frac{3}{2}}\left(x\right)\right)^{2}}{\left(\left(l+1\right)J_{l+\frac{1}{2}}\left(x\right)-xJ_{l+\frac{3}{2}}\left(x\right)\right)^{2}+\left(\left(l+1\right)Y_{l+\frac{1}{2}}\left(x\right)-xY_{l+\frac{3}{2}}\left(x\right)\right)^{2}}$ (11)
where $x\equiv kR$.  For these examples, there are no fields inside the sphere.
For dielectric spheres, on the other hand, the electric and magnetic fields are non-zero within the sphere.  With $\mu=\mu_{0}$ and real, constant index of refraction inside the sphere, $n^{2}=\varepsilon/\varepsilon_{0}$, the standard electromagnetic boundary conditions yield cross-sections that are given by the following partial wave expansions .
$\frac{1}{\pi R^{2}}~\sigma_{\text{dielectric sphere TE + TM}}=\frac{2}{x^{2}}\sum_{l=1}^{\infty}\left(2l+1\right)\left(\left|\frac{S_{l}^{TE}\left(x\right)-1}{2i}\right|^{2}+\left|\frac{S_{l}^{TM}\left(x\right)-1}{2i}\right|^{2}\right)\text{ , \ \ }x\equiv kR$ (12)
$S_{l}^{TE}\left(kR\right)=-\frac{j_{l}\left(nkR\right)\frac{d}{dR}\left(R~h_{l}^{\left(2\right)}\left(kR\right)\right)-h_{l}^{\left(2\right)}\left(kR\right)\frac{d}{dR}\left(R~j_{l}\left(nkR\right)\right)}{j_{l}\left(nkR\right)\frac{d}{dR}\left(R~h_{l}^{\left(1\right)}\left(kR\right)\right)-h_{l}^{\left(1\right)}\left(kR\right)\frac{d}{dR}\left(R~j_{l}\left(nkR\right)\right)}$ (13)
$S_{l}^{TM}\left(kR\right)=-\frac{h_{l}^{\left(2\right)}\left(kR\right)\frac{d}{dR}\left(R~j_{l}\left(nkR\right)\right)-n^{2}j_{l}\left(nkR\right)\frac{d}{dR}\left(R~h_{l}^{\left(2\right)}\left(kR\right)\right)}{h_{l}^{\left(1\right)}\left(kR\right)\frac{d}{dR}\left(R~j_{l}\left(nkR\right)\right)-n^{2}j_{l}\left(nkR\right)\frac{d}{dR}\left(R~h_{l}^{\left(1\right)}\left(kR\right)\right)}$ (14)
For real $n$ the scattering is once again purely elastic with $\left|S_{l}^{TE}\right|=1=\left|S_{l}^{TM}\right|$ for all $k$, and hence $S_{l}^{TE}=\exp\left(2i\delta_{l}\right)$ and $S_{l}^{TM}=\exp\left(2i\Delta_{l}\right)$ with real phase shifts $\delta_{l}$ and $\Delta_{l}$, albeit now given by expressions more complicated than the perfectly conducting case.  For non-resistive material with real $n$, there is no electromagnetic power loss, i.e. no energy absorption within the sphere.  However, energy absorption does take place if the material within the sphere is conductive with resistivity $\rho$.
The simplest model for resistive material is obtained by invoking a linear Ohm’s law relation within the sphere, $\overrightarrow{J}=\overrightarrow{E}/\rho$.  If in addition $\varepsilon\neq\varepsilon_{0}$ but $\mu=\mu_{0}$, then for monochromatic waves
$n^{2}=\frac{1}{\varepsilon_{0}}\left(\varepsilon+\frac{i}{\omega\rho}\right)\ ,\ \ \ \omega=kc$ (15)
The previous results then hold with the substitution $\varepsilon\rightarrow\varepsilon+i/\left(\omega\rho\right)$.  Note the SI units $\left[\rho\right]=[Ohm~Meter]$.  So $\rho/\left(Z_{0}R\right)$ and $\varepsilon_{0}\omega\rho=x\rho/\left(Z_{0}R\right)$ are dimensionless numbers, with $Z_{0}=\sqrt{\mu_{0}/\varepsilon_{0}}\approx 377~Ohms$ and dimensionless $x=kR$.  For this simple model, an obvious method to obtain numerical results approaching those for a perfectly conducting sphere is to take decreasing values for the resistivity $\rho$.
For non-relativistic QM scattering, effects similar to those for non-resistive dielectric spheres can be obtained by considering finite, real, constant potentials within the sphere [9].  By making the potential complex valued, it is also possible to exhibit QM scattering effects similar to those produced by a conductive dielectric sphere with non-zero resistivity.
A more detailed discussion of this QM model is helpful for the remainder of the paper.  Outside the sphere, $V=0$, and the stationary state wave function with energy $\hbar^{2}k^{2}/\left(2m\right)$ is given by the partial wave expansion,
$\psi\left(r,\theta\right)=\sum_{l=0}^{\infty}i^{l}~\frac{1}{2}\left(S_{l}~h_{l}^{\left(1\right)}\left(kr\right)+h_{l}^{\left(2\right)}\left(kr\right)\right)\left(2l+1\right)P_{l}\left(\cos\theta\right)\text{ \ \ for \ \ }R\leq r\leq\infty$ (16)
The scattering amplitude, the differential and integrated elastic cross-sections, and the total cross-section are then given by the general relations,
$f\left(\theta\right)$ $=\frac{1}{k}\sum_{l=0}^{\infty}\left(2l+1\right)\left(\frac{S_{l}-1}{2i}\right)P_{l}\left(\cos\theta\right)\ ,\ \ \ \frac{d\sigma_{el}}{d\Omega}=\left|f\left(\theta\right)\right|^{2}\ ,\ \ \ \sigma_{tot}=\sigma_{el}+\sigma_{inel}$ (17)
[…]
where the inelastic cross-section is defined as $\sigma_{inel}=\sigma_{tot}-\sigma_{el}=\frac{\pi}{k^{2}}\sum_{l=0}^{\infty}\left(2l+1\right)\left(1-\left|S_{l}\right|^{2}\right)$.
Inside the sphere, let
$V\left(r\right)=\frac{\hbar^{2}\eta}{2m}\text{ \ \ for \ \ }0\leq r\leq R$ (19)
Define $x\equiv kR$, $u\equiv\eta R^{2}$, and $\kappa^{2}\equiv k^{2}-\eta$, so $\kappa R=\sqrt{x^{2}-u}$.  Continuity of $\psi$ and it’s first derivatives at $r=R$ leads to
$S_{l}=-\left(\frac{x~J_{l+1/2}\left(\sqrt{x^{2}-u}\right)H_{l+3/2}^{\left(2\right)}\left(x\right)-\sqrt{x^{2}-u}~J_{l+3/2}\left(\sqrt{x^{2}-u}\right)H_{l+1/2}^{\left(2\right)}\left(x\right)}{x~J_{l+1/2}\left(\sqrt{x^{2}-u}\right)H_{l+3/2}^{\left(1\right)}\left(x\right)-\sqrt{x^{2}-u}~J_{l+3/2}\left(\sqrt{x^{2}-u}\right)H_{l+1/2}^{\left(1\right)}\left(x\right)}\right)$ (20)
The coefficients appearing in $f\left(\theta\right)$ and $\sigma_{el}$ are then
$A_{l}=\frac{i}{2}\left(1-S_{l}\right)=i\left(\frac{x~J_{l+1/2}\left(\sqrt{x^{2}-u}\right)J_{l+3/2}\left(x\right)-\sqrt{x^{2}-u}~J_{l+3/2}\left(\sqrt{x^{2}-u}\right)J_{l+1/2}\left(x\right)}{x~J_{l+1/2}\left(\sqrt{x^{2}-u}\right)H_{l+3/2}^{\left(1\right)}\left(x\right)-\sqrt{x^{2}-u}~J_{l+3/2}\left(\sqrt{x^{2}-u}\right)H_{l+1/2}^{\left(1\right)}\left(x\right)}\right)$ (21)
For real $\eta$ (i.e. real $u=\eta R^{2}$) it follows that $\left|S_{l}\right|=1$ for all real $k$ (i.e. real $x=kR$).  In this case $\sigma_{inel}=0$.  But in general, for complex $\eta$ (i.e. complex $u$), $\left|S_{l}\right|\neq 1$ and $\sigma_{inel}\neq 0$.
3 Selected spatial geometries
A simple “foxhole” geometric model that is able to mimic features of Mie scattering for real index of refraction is given by endowing a three dimensional manifold with a non-trivial metric as obtained by the following $\mathbb{M}_{3}\subset\mathbb{E}_{4}$ embedding.
$\left(ds\right)^{2}=\left(dh\right)^{2}+\left(dr\right)^{2}+r^{2}\left(d\theta\right)^{2}+\left(r^{2}\sin^{2}\theta\right)\left(d\phi\right)^{2}$ (22)
[…]
where $\theta$ and $\phi$ are the usual spherical polar angles in $\mathbb{E}_{3}$.  For example, when $p=1/2$, equatorial slices of the manifold for any fixed $\phi$, for various $n$, are pictured here:
[…]
Analytically, the limit $n\rightarrow\infty$ is somewhat obscure, but from the previous graph, that limit is clear.  As $n\rightarrow\infty$ the geometry is a cylinder of height $H$ composed of 2-spheres each of radius $R$ (i.e. $\left[-H,0\right]\otimes S_{2}\left(R\right)$) terminated at the “lower” end by a ball of radius $R$ (i.e. $\mathbb{B}_{3}\left(R\right)$) and attached at the “upper” end to a punctured $\mathbb{E}_{3}$ (i.e. $\mathbb{E}_{3}\smallsetminus\mathbb{B}_{3}\left(R\right)$).  It will turn out that scattering from this foxhole geometry is elastic for real $H$, but if $H$ is taken to be complex, the model can produce inelastic scattering.
Alternatively, a geometric model obtained from a wormhole metric also exhibits features of Mie scattering with absorption.  Consider a smooth, spatial “bridge” manifold [12] (i.e. a static “wormhole” [13]) defined by
$\left(ds\right)^{2}$ $=R^{2}\left(dw\right)^{2}+r^{2}\left(w\right)\left(d\theta\right)^{2}+\left(r^{2}\left(w\right)\sin^{2}\theta\right)\left(d\phi\right)^{2}$ (24)
$r\left(w\right)$ $=R\left(1+\left(w^{2}\right)^{n/2}\right)^{1/n}\ ,\ \ \ -\infty\leq w\leq+\infty$ (25)
For example, the static version of the Ellis metric [14] is given by $n=2$, namely,
$r\left(w\right)=R\sqrt{1+w^{2}}\ ,\ \ \ -\infty\leq w\leq+\infty$
$x\left(w,\theta\right)=r\left(w\right)\cos\theta\ ,\ \ \ y\left(w,\theta\right)=r\left(w\right)\sin\theta$ (26)
As before, a 2D equatorial slice of the manifold (i.e. $\theta=\pi/2$, hence $z=0$) can be embedded in 3D, only now with the help of an “extra-physical” dimension, $h$.  On that slice
$\left(ds\right)^{2}=R^{2}\left(dw\right)^{2}+r^{2}\left(w\right)\left(d\phi\right)^{2}=\left(dx\right)^{2}+\left(dy\right)^{2}+\frac{1}{\alpha^{2}}\left(dh\right)^{2}$ (27)
$x\left(w,\phi\right)=r\left(w\right)\cos\phi\ ,\ \ \ y\left(w,\phi\right)=r\left(w\right)\sin\phi$ (28)
[…]
where $\alpha\equiv H/R$ is the “aspect ratio” of the wormhole.  For example, for the Ellis case with $\alpha=1$ and $n=2$,
$h\left(w\right)=R\ln\left(w+\sqrt{1+w^{2}}\right)=R\operatorname{arcsinh}\left(w\right)$ (30)
For generic $n$, it is easiest to obtain $h\left(w\right)$ by numerical solution of
$\frac{1}{R}\frac{dh\left(w\right)}{dw}=\alpha\sqrt{1-\left(w^{2}\right)^{n-1}\left(1+\left(w^{2}\right)^{n/2}\right)^{\frac{2}{n}-2}}$ (31)
with initial condition $h\left(0\right)=0$.
Here are equatorial slice profiles for the $\left(x,h\right)$ plane of the embedded manifolds, for five different values of $n$.  The same profile is obtained for the $\left(y,h\right)$ plane, or for any other plane oriented at a fixed azimuthal angle.  That is to say, the embedding is rotationally invariant about the $h$ axis.
[…]
It so happens that scattering from this wormhole geometry is inelastic for any $n\geq 2$, in the sense that the wormhole absorbs particle flux incident upon it from infinite distance, say, on the upper branch of the manifold.  In the rest of this paper we will consider mostly the $n=\infty$ foxhole and wormhole geometries, with just a few related remarks for the $n=2$ Ellis wormhole.
Variants of these foxhole and wormhole geometries are easily imagined.  One such variant would be to join the finite $n$ geometries smoothly with flat space at a fixed radius $R_{1}>R$.  For example, with Heaviside step function $\Theta$, a modified Ellis wormhole is given by
$r\left(w\right)=R\left(1+\left(w^{2}\right)^{n/2}\right)^{1/n}\Theta\left(w_{1}^{2}-w^{2}\right)+R_{1}\sqrt{w^{2}/w_{1}^{2}}~\Theta\left(w^{2}-w_{1}^{2}\right),\ \ \ -\infty\leq w\leq+\infty$ (32)
where $R_{1}=R\left(1+\left(w_{1}^{2}\right)^{n/2}\right)^{1/n}$ for some chosen $w_{1}>0$.  This modification would give a continuous metric whose first derivatives are discontinuous at $R_{1}$, and whose second derivatives are infinite (i.e. Dirac delta terms in the curvature) at $R_{1}$.  (This is also true for either of the $n\rightarrow\infty$ manifolds at $r=R$.)  However, for the models to be considered the wave function $\Psi$ is taken to be a solution to the covariant Helmholtz equation,
$\left(\nabla^{2}+k^{2}\right)\Psi=0$ (33)
where $\nabla^{2}$ is the invariant Laplacian on the manifold, $\nabla^{2}\Psi=\frac{1}{\sqrt{g}}\partial_{\mu}\left(\sqrt{g}g^{\mu\nu}\partial_{\nu}\Psi\right)$.  Therefore, without coupling $\Psi$ directly to the curvature, $\Psi$ and its first derivatives are allowed to be continuous at $R=R_{1}$, so there will be no Dirac deltas in this covariant Helmholtz equation.  (This is also true for either of the $n\rightarrow\infty$ manifolds at $r=R$.)
4 Amplitudes
The $n\rightarrow\infty$ foxhole geometry consists of three flat space parts:  The punctured space $\mathbb{E}_{3}\smallsetminus\mathbb{B}_{3}\left(R\right)$, the cylinder composed of $S_{2}\left(R\right)$s, and the ball $\mathbb{B}_{3}\left(R\right)$ at $h=-H$.  On $\mathbb{E}_{3}\smallsetminus\mathbb{B}_{3}\left(R\right)$ and on $\mathbb{B}_{3}\left(R\right)$,
$\nabla^{2}\Psi=\frac{1}{r}\frac{\partial^{2}}{\partial r^{2}}\left(r\Psi\right)-\frac{1}{r^{2}}~L^{2}\Psi$ (34)
where $\overrightarrow{L}=-i\overrightarrow{r}\times\overrightarrow{\nabla}$.  Upon choosing angular momentum eigenstates, and factoring $\Psi=\psi\left(r\right)Y_{lm}\left(\theta,\phi\right)$, the radial function must satisfy
$\frac{1}{r}\frac{\partial^{2}}{\partial r^{2}}\left(r\psi\right)+\left(k^{2}-\frac{l\left(l+1\right)}{r^{2}}\right)\psi=0$ (35)
with spherical Bessel function solutions $\psi=j_{l}\left(kr\right)=\sqrt{\frac{\pi}{2kr}}J_{l+1/2}\left(kr\right)$ or $\psi=y_{l}\left(kr\right)=\sqrt{\frac{\pi}{2kr}}Y_{l+1/2}\left(kr\right)$.  On the cylinder
$\nabla^{2}\Psi=\frac{\partial^{2}}{\partial h^{2}}\Psi-\frac{1}{R^{2}}~L^{2}\Psi$ (36)
with definite angular momentum states given by $\Psi=\psi\left(h\right)Y_{lm}\left(\theta,\phi\right)$. Helmholtz’s equation is now satisfied provided
$\frac{\partial^{2}}{\partial h^{2}}\psi\left(h\right)+\left(k^{2}-\frac{l\left(l+1\right)}{R^{2}}\right)\psi\left(h\right)$ (37)
with solutions $\psi=\sin\left(\kappa h\right)$ and $\psi=\cos\left(\kappa h\right)$ where $\kappa^{2}=k^{2}-l\left(l+1\right)/R^{2}$.
For an incident plane wave $\exp\left(ikr\cos\theta\right)$ only the $\phi$-independent $P_{l}\left(\cos\theta\right)=\sqrt{\frac{4\pi}{2l+1}}Y_{l0}\left(\theta,\phi\right)$ are involved in the scattering solutions.  In this case the solutions on $\mathbb{E}_{3}\smallsetminus\mathbb{B}_{3}\left(R\right)$ take the familiar form
$\Psi\left(r,\theta\right)=\sum_{l=0}^{\infty}i^{l}~\frac{1}{2}\left(S_{l}~h_{l}^{\left(1\right)}\left(kr\right)+h_{l}^{\left(2\right)}\left(kr\right)\right)\left(2l+1\right)P_{l}\left(\cos\theta\right)\text{ \ \ for \ \ }R<r\leq\infty$ (38)
with coefficients $S_{l}$ to be determined, while on the ball
$\Psi\left(r,\theta\right)=\sum_{l=0}^{\infty}i^{l}~\frac{1}{2}B_{l}~j_{l}\left(kr\right)\left(2l+1\right)P_{l}\left(\cos\theta\right)\text{ \ \ for \ \ }0\leq R<r$ (39)
since the solution must be regular at $r=0$.  The solutions on the cylinder take the form
$\Psi\left(h,\theta\right)=\sum_{l=0}^{\infty}i^{l}~\frac{1}{2}\left(C_{l}~\cos\left(\kappa h\right)+D_{l}~\sin\left(\kappa h\right)\right)\left(2l+1\right)P_{l}\left(\cos\theta\right)$ (40)
Continuity of $\Psi$ and its first derivatives on the edges of the cylinder now determines all the coefficients.
That is to say, the wave function must satisfy the boundary conditions
$\left.\Psi\right|_{h=0,\ r\searrow R}$ $=\left.\Psi\right|_{r=R,\ h\nearrow 0}\text{ \ \ and \ \ }\left.\frac{\partial}{\partial r}\Psi\right|_{h=0,\ r\searrow R}=\left.\frac{\partial}{\partial h}\Psi\right|_{r=R,\ h\nearrow 0}$ (41)
$\left.\Psi\right|_{h=-H,\ r\nearrow R}$ $=\left.\Psi\right|_{r=R,\ h\searrow-H}\text{ \ \ and \ \ }\left.\frac{\partial}{\partial r}\Psi\right|_{h=-H,\ r\nearrow R}=\left.\frac{\partial}{\partial h}\Psi\right|_{r=R,\ h\searrow-H}$ (42)
Upon using the expansions for $\Psi$ on the three parts of the $n\rightarrow\infty$ manifold, the coefficients are determined by these boundary conditions.  The results are
$S_{l}\left(x,y\right)$ $=-\frac{\operatorname*{numer}\left(l,x,y\right)}{\operatorname*{denom}\left(l,x,y\right)}\ ,\ \ \ B_{l}\left(x,y\right)=\frac{4i}{\pi}\frac{\sqrt{x^{2}-l\left(l+1\right)}}{\operatorname*{denom}\left(l,x,y\right)}\ ,$ (43)
$C_{l}\left(x,y\right)$ $=+i\sqrt{\frac{8}{\pi x}}\frac{\sqrt{x^{2}-l\left(l+1\right)}J_{l+1/2}\left(x\right)\cos\left(2\sqrt{x^{2}-l\left(l+1\right)}y\right)}{\operatorname*{denom}\left(l,x,y\right)}\ ,$ (44)
$D_{l}\left(x,y\right)$ $=-i\sqrt{\frac{8}{\pi x}}\frac{\sqrt{x^{2}-l\left(l+1\right)}J_{l+1/2}\left(x\right)\sin\left(2y\sqrt{x^{2}-l\left(l+1\right)}\right)}{\operatorname*{denom}\left(l,x,y\right)}\ ,$ (45)
with the definitions (not to be confused with Cartesian coordinates for $\mathbb{E}_{3}$!)
$x\equiv kR\ ,\ \ \ y\equiv H/\left(2R\right)\ ,\ \ \ \kappa R=\sqrt{x^{2}-l\left(l+1\right)}\equiv z\left(l,x\right)\ ,\ \ \ \kappa H/2=y\sqrt{x^{2}-l\left(l+1\right)}=y~z\left(l,x\right)$ (46)
and where the denominator and numerator functions are given by
$\operatorname*{denom}\left(l,x,y\right)$ $=+\frac{2}{\pi}i\sqrt{x^{2}-l\left(l+1\right)}\cos\left(2y~z\left(l,x\right)\right)$ (47)
$+\left(\left(\left(x^{2}-l\right)J_{l+\frac{1}{2}}\left(x\right)-lxJ_{l+\frac{3}{2}}\left(x\right)\right)H_{l+\frac{1}{2}}^{\left(1\right)}\left(x\right)+x\left(xJ_{l+\frac{3}{2}}\left(x\right)-lJ_{l+\frac{1}{2}}\left(x\right)\right)H_{l+\frac{3}{2}}^{\left(1\right)}\left(x\right)\right)\sin\left(2y~z\left(l,x\right)\right)$
$\operatorname*{numer}\left(l,x,y\right)$ $=-\frac{2}{\pi}i\sqrt{x^{2}-l\left(l+1\right)}\cos\left(2y~z\left(l,x\right)\right)$ (48)
$+\left(\left(\left(x^{2}-l\right)J_{l+\frac{1}{2}}\left(x\right)-lxJ_{l+\frac{3}{2}}\left(x\right)\right)H_{l+\frac{1}{2}}^{\left(2\right)}\left(x\right)+x\left(xJ_{l+\frac{3}{2}}\left(x\right)-lJ_{l+\frac{1}{2}}\left(x\right)\right)H_{l+\frac{3}{2}}^{\left(2\right)}\left(x\right)\right)\sin\left(2y~z\left(l,x\right)\right)$
For real $x$ and real $y$ note that $\operatorname*{numer}\left(l,x,y\right)=\operatorname*{denom}^{\ast}\left(l,x,y\right)$ if $x^{2}\geq l\left(l+1\right)$ and $\operatorname*{numer}\left(l,x,y\right)=-\operatorname*{denom}^{\ast}\left(l,x,y\right)$ if $x^{2}\leq l\left(l+1\right)$.  But either way, no matter whether $x^{2}\geq l\left(l+1\right)$ or $x^{2}\leq l\left(l+1\right)$, the scattering is elastic: $\left|S_{l}\left(x,y\right)\right|=1$.  The coefficients appearing in $f\left(\theta\right)$ and $\sigma_{el}$ are then
$A_{l}$ $=\frac{i}{2}\left(1-S_{l}\right)$ (49)
$=i\frac{\left(\left(\left(x^{2}-l\right)J_{l+\frac{1}{2}}\left(x\right)-lxJ_{l+\frac{3}{2}}\left(x\right)\right)J_{l+\frac{1}{2}}\left(x\right)+x\left(xJ_{l+\frac{3}{2}}\left(x\right)-lJ_{l+\frac{1}{2}}\left(x\right)\right)J_{l+\frac{3}{2}}\left(x\right)\right)\sin\left(2y~z\left(l,x\right)\right)}{\operatorname*{denom}\left(l,x,y\right)}$
For the $n\rightarrow\infty$ wormhole geometry, with incident flux only on the “upper” copy of $\mathbb{E}_{3}\smallsetminus\mathbb{B}_{3}\left(R\right)$, the difference with the foxhole partial wave expansions is just the existence of purely outgoing wave Bessel functions (i.e. $h_{l}^{\left(1\right)}\left(kr\right)$ but not $h_{l}^{\left(2\right)}\left(kr\right)$) on the “lower” copy of $\mathbb{E}_{3}\smallsetminus\mathbb{B}_{3}\left(R\right)$ instead of the regular Bessel functions (i.e. $j_{l}\left(kr\right)$) on the foxhole ball $\mathbb{B}_{3}\left(R\right)$.  On that lower $\mathbb{E}_{3}\smallsetminus\mathbb{B}_{3}\left(R\right)$ the solution is now given by the expansion
$\Psi\left(r,\theta\right)=\sum_{l=0}^{\infty}i^{l}~\frac{1}{2}~T_{l}~h_{l}^{\left(1\right)}\left(kr\right)\left(2l+1\right)P_{l}\left(\cos\theta\right)\text{ \ \ for \ \ }R<r\leq\infty$ (50)
with coefficients $T_{l}$ to be determined along with $S_{l}$, $C_{l}$, and $D_{l}$, where the latter are defined as in the foxhole expansions.  Corresponding expansions apply if there is incident flux only on the lower 3D space, of course, but only incident flux on the upper 3D space is considered here.
The wave function must now satisfy the boundary conditions
$\left.\Psi\right|_{h=H,\ r\searrow R}$ $=\left.\Psi\right|_{r=R,\ h\nearrow H}\text{ \ \ and \ \ }\left.\frac{\partial}{\partial r}\Psi\right|_{h=H,\ r\searrow R}=\left.\frac{\partial}{\partial h}\Psi\right|_{r=R,\ h\nearrow H}$ (51)
$\left.\Psi\right|_{h=-H,\ r\searrow R}$ $=\left.\Psi\right|_{r=R,\ h\searrow-H}\text{ \ \ and \ \ }\left.\frac{\partial}{\partial r}\Psi\right|_{h=-H,\ r\searrow R}=-\left.\frac{\partial}{\partial h}\Psi\right|_{r=R,\ h\searrow-H}$ (52)
with a crucial minus sign in the last of these conditions.  Upon using the expansions for $\Psi$ on the three parts of the $n\rightarrow\infty$ manifold, all the coefficients are determined by these boundary conditions.  The results most relevant for scattering are
$S_{l}\left(x,y\right)=-\frac{\operatorname*{wumer}\left(l,x,y\right)}{\operatorname*{wenom}\left(l,x,y\right)}\ ,\ \ \ T_{l}\left(x,y\right)=\frac{4i}{\pi}\frac{\sqrt{x^{2}-l\left(l+1\right)}}{\operatorname*{wenom}\left(l,x,y\right)}$ (53)
where the wormhole denominator and numerator functions are given by
$\operatorname*{wenom}\left(l,x,y\right)$ $=2\sqrt{x^{2}-l\left(l+1\right)}H_{l+1/2}^{\left(1\right)}\left(x\right)\left(lH_{l+1/2}^{\left(1\right)}\left(x\right)-xH_{l+3/2}^{\left(1\right)}\left(x\right)\right)\cos\left(2y~z\left(l,x\right)\right)$ (54)
$+\left(\left(x^{2}-l\left(l+1\right)\right)\left(H_{l+1/2}^{\left(1\right)}\left(x\right)\right)^{2}-\left(lH_{l+1/2}^{\left(1\right)}\left(x\right)-xH_{l+3/2}^{\left(1\right)}\left(x\right)\right)^{2}\right)\sin\left(2y~z\left(l,x\right)\right)$
$\operatorname*{wumer}\left(l,x,y\right)=$ (55)
$\sqrt{x^{2}-l\left(l+1\right)}\left(H_{l+1/2}^{\left(2\right)}\left(x\right)\left(lH_{l+1/2}^{\left(1\right)}\left(x\right)-xH_{l+3/2}^{\left(1\right)}\left(x\right)\right)+H_{l+1/2}^{\left(1\right)}\left(x\right)\left(lH_{l+1/2}^{\left(2\right)}\left(x\right)-xH_{l+3/2}^{\left(2\right)}\left(x\right)\right)\right)\cos\left(2y~z\left(l,x\right)\right)$
$+\left(\left(x^{2}-l\left(l+1\right)\right)H_{l+1/2}^{\left(2\right)}\left(x\right)H_{l+1/2}^{\left(1\right)}\left(x\right)-\left(lH_{l+1/2}^{\left(2\right)}\left(x\right)-xH_{l+3/2}^{\left(2\right)}\left(x\right)\right)\left(lH_{l+1/2}^{\left(1\right)}\left(x\right)-xH_{l+3/2}^{\left(1\right)}\left(x\right)\right)\right)\sin\left(2y~z\left(l,x\right)\right)$
with the definitions (again, not to be confused with Cartesian coordinates for $\mathbb{E}_{3}$!)
$x\equiv kR\ ,\ \ \ y\equiv H/R\ ,\ \ \ \kappa R=\sqrt{x^{2}-l\left(l+1\right)}\equiv z\left(l,x\right)\ ,\ \ \ \kappa H=y\sqrt{x^{2}-l\left(l+1\right)}=y~z\left(l,x\right)$ (56)
Note the definition of the aspect ratio $y$ in terms of $H$ and $R$ for the wormhole differs from that for the foxhole by a factor of 222.
For the wormhole with incident waves only on the upper 3D space, it does not matter if $k^{2}R^{2}\geq l\left(l+1\right)$ so that $\kappa R=\sqrt{k^{2}R^{2}-l\left(l+1\right)}$ is real, or if $k^{2}R^{2}\leq l\left(l+1\right)$ so that $\kappa R=i\sqrt{l\left(l+1\right)-k^{2}R^{2}}$ is imaginary, either way $\left|S_{l}\right|\neq 1$.  This means there is not only elastic scattering on the upper 3D space but also absorption by the wormhole with outward flow on the bottom 3D space.  The elastic cross section for scattered waves on the upper 3D space is given by
$\sigma_{el}=\frac{\pi}{k^{2}}\sum_{l=0}^{\infty}\left(2l+1\right)\left|S_{l}-1\right|^{2}=\frac{\pi}{k^{2}}\sum_{l=0}^{\infty}\left(2l+1\right)\left(1+\left|S_{l}\right|^{2}-2\operatorname{Re}S_{l}\right)$ (57)
while the inelastic (absorption) and total cross-sections are given by
$\sigma_{inel}=\frac{\pi}{k^{2}}\sum_{l=0}^{\infty}\left(2l+1\right)\left(1-\left|S_{l}\right|^{2}\right)\ ,\ \ \ \sigma_{tot}=\sigma_{el}+\sigma_{inel}=\frac{2\pi}{k^{2}}\sum_{l=0}^{\infty}\left(2l+1\right)\left(1-\operatorname{Re}S_{l}\right)$ (58)
Since the model is invariant under $\left.\mathbb{E}_{3}\right|_{\text{upper}}\leftrightarrow\left.\mathbb{E}_{3}\right|_{\text{lower}}$, if there is incident flux only on the lower 3D space, with corresponding expansion coefficients, the same elastic and inelastic cross-sections are obtained.
5 Resonance poles
The foxhole scattering amplitudes exhibit poles in the lower half complex $k$ (or $x\equiv kR$) plane, in complete accord with general theory for potential scattering [9]. For the $n\rightarrow\infty$ foxhole geometry these poles can be obtained to arbitrary accuracy by numerical computation of the $\operatorname*{denom}\left(l,x,y\right)$ zeroes. Alternatively, a simple estimate of the corresponding resonance frequencies follows from the Bohr-Sommerfeld (B-S) method [15].
The cylinder is all that distinguishes the $n\rightarrow\infty$ foxhole geometry from flat space.  On the cylinder the classical action for a complete “up and down” cycle of motion along the cylinder length is given by a trivial integral, since momentum is constant on the cylinder.
[…]
with $x=kR$ and $y=H/\left(2R\right)$.  Quantizing this action as $I=\left(2n+1\right)\pi$ for $n=0,1,2,\cdots$ then gives an estimate of the various resonant $x$ values as
$x_{res}=\sqrt{l\left(l+1\right)+\frac{\left(2n+1\right)^{2}\pi^{2}}{16y^{2}}}$ (60)
[…]
$x_{l=0\text{ resonance}}=\pi/8\ ,\ \ \ 3\pi/8\ ,\ \ \ 5\pi/8,\ \ \ 7\pi/8\ ,\ \ \ \cdots\text{ \ \ }=0.3927\ ,\ \ \ 1.178\ ,\ \ \ 1.963\ ,\ \ \ 2.749\ ,\ \ \ \cdots$ (61)
The accuracy of these estimates for $l=0$ leaves something to be desired, as evident in the numerical solution of $0=\operatorname*{denom}\left(0,x,2\right)$ as well as numerical plots of the partial cross-section $\sigma_{0}\left(x,2\right)$.
However, the B-S estimate for the cross-section peaks improves as $kR$ becomes large.  To understand this, consider the principal asymptotic forms for the Bessel functions that make up the scattering amplitudes.
$J_{l+\frac{1}{2}}\left(x\right)\underset{x\gg l}{\sim}\sqrt{\frac{2}{\pi x}}\sin\left(x-\frac{\pi}{2}l\right)\ ,\ \ \ Y_{l+\frac{1}{2}}\left(x\right)\underset{x\gg l}{\sim}-\sqrt{\frac{2}{\pi x}}\cos\left(x-\frac{\pi}{2}l\right)$ (62)
$H_{l+\frac{1}{2}}^{\left(1\right)}\left(x\right)\underset{x\gg l}{\sim}-i\sqrt{\frac{2}{\pi x}}e^{i\left(x-\frac{\pi}{2}l\right)}\ ,\ \ \ H_{l+\frac{1}{2}}^{\left(2\right)}\left(x\right)\underset{x\gg l}{\sim}i\sqrt{\frac{2}{\pi x}}e^{-i\left(x-\frac{\pi}{2}l\right)}$ (63)
It follows that the leading asymptotic behavior of $\operatorname*{denom}\left(l,x,y\right)$ is
$\operatorname*{denom}\left(l,x,y\right)\underset{x\gg l}{\sim}\frac{2ix}{\pi}\exp\left(-2iy\sqrt{x^{2}-l\left(l+1\right)}\right)$ (64)
with real and imaginary parts
$\operatorname{Re}\left(\operatorname*{denom}\left(l,x,y\right)\right)\underset{x\gg l}{\sim}\frac{2x}{\pi}~\sin\left(2y\sqrt{x^{2}-l\left(l+1\right)}\right)$ (65)
$\operatorname{Im}\left(\operatorname*{denom}\left(l,x,y\right)\right)\underset{x\gg l}{\sim}\frac{2x}{\pi}~\cos\left(2y\sqrt{x^{2}-l\left(l+1\right)}\right)$ (66)
Roots of the asymptotic form for $\operatorname{Im}\left(\operatorname*{denom}\left(l,x,y\right)\right)$ then yield the B-S quantization condition:  $2y\sqrt{x^{2}-l\left(l+1\right)}=\left(n+1/2\right)\pi$.
On the other hand, evaluating the asymptotic form of $\operatorname{Re}\left(\operatorname*{denom}\left(l,x,y\right)\right)$ at these roots then gives
$\left.\operatorname{Re}\left(\operatorname*{denom}\left(l,x,y\right)\right)\right|_{\text{B-S}}\underset{x\gg l}{\sim}\frac{2x}{\pi}\left(-1\right)^{n}$ (67)
which does not vanish.  But in fact, this is just what should have been expected, since the asymptotic form of the scattering amplitude itself is
$S_{l}\left(x,y\right)\underset{x\gg l}{\sim}\exp\left(4iy\sqrt{x^{2}-l\left(l+1\right)}\right)=\exp\left(2iH\sqrt{k^{2}-\frac{l\left(l+1\right)}{R^{2}}}\right)$ (68)
where once again $y=\frac{H}{2R}$.  This is just the additional phase acquired by a plane wave with momentum $\kappa=\sqrt{k^{2}-\frac{l\left(l+1\right)}{R^{2}}}$ as it goes down the cylinder length, and then back up the cylinder length, with no other net phase change from its encounters with both the upper and lower edges of the cylinder.
Evaluating at the B-S condition gives
$\left.S_{l}\left(x,y\right)\right|_{\text{B-S}}\underset{x\gg l}{\sim}\exp\left(\left(2n+1\right)i\pi\right)=-1$ (69)
This value for $S_{l}$ maximizes $\left|A_{l}\right|$ and the partial cross-section.  To be more explicit, the asymptotic form of the $l$th partial wave total cross-section for scattering from the foxhole is given by
$\frac{\sigma_{l}\left(x,y\right)}{\pi R^{2}}\underset{x\gg l}{\sim}\frac{2}{x^{2}}\left(2l+1\right)\left(1-\operatorname{Re}\left(\exp\left(4iy\sqrt{x^{2}-l\left(l+1\right)}\right)\right)\right)=\frac{4}{x^{2}}\left(2l+1\right)\sin^{2}\left(2y\sqrt{x^{2}-l\left(l+1\right)}\right)$ (70)
Evaluating at the B-S condition then saturates the unitarity bound for this partial cross-section:
$\left.\frac{\sigma_{l}\left(x,y\right)}{\pi R^{2}}\right|_{\text{B-S}}\underset{x\gg l}{\sim}\frac{4}{x^{2}}\left(2l+1\right)$ (71)
The asymptotic partial cross-section is therefore maximized at the B-S value for $kR$.
6 Cross-sections
Consider the $n\rightarrow\infty$ foxhole with aspect ratio $y=2$, and compute numerically the net elastic cross-section for $x\leq 6$.  This is given accurately by $\sigma_{el}=\sum_{l=0}^{5}\sigma_{l}\left(x,2\right)$ since for this aspect ratio the higher $l$ partial waves only contribute significantly for $x>6$.  The numerical results are as shown here.
Figure 3: Fox Hole Scattering
Figure 4: Fox Hole Scattering
Note the change in vertical scales in these two graphs.  Also recall that $\sigma_{tot}=\sigma_{el}$ for real aspect ratios.
As happens in the case of Mie scattering for real index of refraction, the complicated structure exhibited by the foxhole total cross-section can be better understood by considering the individual partial wave contributions.  The first few of these partial cross-sections are plotted below.  The light gray curves are unitarity bounds, and the light blue lines indicate B-S approximations for the real parts of the resonance poles located in the lower half of the complex $k$ plane, as discussed in the previous Section.  For example, consider $l=0$.
Figure 5: Resonance for $\leq$0.5
[…]
Figure 7: for l=1
Figure 8: for l=2
Figure 9: for l=3
The accuracy of the approximation $\sigma_{el}=\sum_{l=0}^{5}\sigma_{l}\left(x,2\right)$ for $x\leq 6$ is evident in the following graph of $\sigma_{6}\left(x,2\right)$.
Figure 10: for l=6
Numerical results for the infinite $n$ wormhole geometry defined above are easily obtained from the exact partial wave expansions of the previous sections, evincing qualitative features similar to those of the foxhole example, but with the added feature of non-zero inelastic scattering (some flux goes “down the drain”) [3, 4].
7 Tractable Generalizations
There are several generalizations of the geometries given here that warrant further study.  For either the foxhole or wormhole models, a tractable, easily solvable modification of the infinite $n$ geometry in Section 3 would be to have different radii for the spheres at the top and bottom of the cylindrical section, i.e. the corresponding equatorial slice of the cylinder would be a trapezoid.  Such modifcations would perhaps convey intuition about related dielectric systems.  If orthogonally projected onto extensions of the ambient Euclidean space (say, as viewed “from above” in diagrams similar to those in Figures 1 and 2), the resulting projection of wavelets on the trapezoidal cylinder would appear to have shorter wavelengths, hence lower phase velocities, thereby corresponding to effective refraction indices with magnitude greater than unity.  Moreover, if the lower radius is greater than the upper radius, the effective projected phase velocity of wavelets on the cylindrical portion of the trapezoidal foxhole would appear to be negative.
8 Summary
Scattering due to nontrivial spherically symmetric spatial geometries, and due to nothing else, has been illustrated in this paper for a selection of simple geometries.  The analysis was non-relativistic in the sense that it was framed in the mathematical context of the three-dimensional Helmholtz equation where only spatial geometry was modified from Euclidean space, with time taken to be universal. The emphasis here was on the geometry and the resulting cross-sections as computed directly from exact results for partial wave expansions.
Relevant formulae were given in the body of the paper along with selected numerical graphs to illustrate various important physical effects.  Among those effects were the occurrence of resonances and the saturation of unitarity bounds.
Additional plots for comparison to various well-known situations have been collected in an Appendix.
Acknowledgements TC received financial support from the United States Social Security Administration.
Disclosures The authors declare no conflicts of interest.
Appendix A Appendix:  Perfect spheres
Figure A-1: In black, individual $\sigma_{l}/\left(\pi R^{2}\right)$ for $l=0,1,2$ 333 from left to right.
In red, $\sigma_{\text{QM hard sphere}}/\left(\pi R^{2}\right)=\sum_{l=0}^{\infty}\sigma_{l\text{ QM hard sphere}}/\left(\pi R^{2}\right)$
but only $l\leq 5$ contribute significantly for $kR\leq 5$.
Figure A-2: $\frac{1}{\pi R^{2}}\left(\sigma_{\text{TE perf cond sphere}}+\sigma_{\text{TM perf cond sphere}}\right)$ and
$\frac{1}{\pi R^{2}}~\sigma_{\text{quasi-static}}=\frac{10}{3}k^{4}R^{4}$ (light blue).
Figure A-3: $\frac{1}{\pi R^{2}}~\sigma_{\text{TE perf cond sphere}}$ and $\frac{2}{3}k^{4}R^{4}$ (light blue)
Figure A-4: $\frac{1}{\pi R^{2}}~\sigma_{\text{TE perf cond sphere}}$ and $\frac{2}{3}k^{4}R^{4}$ (light blue)
For dielectric spheres, consider the case $n=2$ for numerical purposes.
Figure A-5: $l=1$ to 666 summed contributions to $\sigma_{TE}$
Figure A-6: $l=1$ to 666 summed contributions to $\sigma_{TM}$.
Also, for real $n$, the so-called “physical optics” approximation for Mie scattering is [11]
$\sigma_{P}\left(x,n\right)=2\left(1-\frac{2}{2x\left(n-1\right)}\sin\left(2x\left(n-1\right)\right)+\frac{2}{\left(2x\left(n-1\right)\right)^{2}}\left(1-\cos\left(2x\left(n-1\right)\right)\right)\right)$
Figure A-7: $l=1$ to 666 summed contributions to $\sigma_{TE}+\sigma_{TM}$
with $\sigma_{P}$ in green.
Figure A-8: $l=1$ to 121212 summed contributions to $\sigma_{TE}+\sigma_{TM}$ in red and $\sigma_{P}$ in green ($l=1$ to 666 summed contributions in orange).